# Advancing Reference-free Evaluation of Video Captions with Factual Analysis

### v1

@@ 4.2 Training
where we format the explanation $E$ in text using the information collected in the data creation process, i.e., the list of changed objects and actions, detailed in the supplementary. This can serve as extra supervision for the model to implicitly learn factual grounding and provide interpretable reasoning for this model-based evaluator at test time. During training, the model is optimized with the language modeling loss, like in other instruction tuning work [20].
@@ 5.1 Experimental Settings
We train VC-Inspector with the proposed instruction tuning dataset, ActivityNet-FG-It, which comprises 44K video-caption pairs, along with their quality scores and explanations. To evaluate whether VC-Inspector can generalize to other videos and visual domains, we further follow the same data generation pipeline in section 4.1 to create two evaluation datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, according to the ActivityNet [14] test set and YouCook2 [41] validation set. However, these are not truly labeled by humans. To evaluate whether the quality scores given by VC-Inspector are aligned with human evaluators, we utilized the widely adopted VATEX-EVAL dataset, specifically designed for evaluating video caption evaluation quality. It contains 6 captions with varying levels of quality per video, each rated by three human evaluators on a scale of 1 to 5. Since some videos become unavailable on YouTube, we collect the remaining subset of 2,590 videos and the corresponding 15,540 candidate captions. Unless otherwise specified, all experiments were evaluated on the same dataset to ensure a fair comparison. We further extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], by viewing images as single-frame videos to test the generalization of the proposed metric. The former contains 17K image-caption pairs rated by three human experts on a scale of 1 to 4, while the latter collects binary quality assessments for 48K image-caption pairs from crowd sources.
@@
The baselines are organized into three categories: i) Language-based metrics: The evaluation in this category solely relies on the reference texts without considering the visual input. Representative metrics are BLEU [26], ROUGE [19], METEOR [3], CIDEr [33], BERTScore [39], SPICE [1], SPICE-Factual [18], Soft-SPICE  [18] and CLAIR [6]. ii) Image-augmented metrics: These approaches incorporate images as references (alongside reference captions), for better capturing the semantic alignment between the visual input and the candidate caption, e.g., CLIPScore [9], EMScore [31], PAC-S [30], FactVC [21], and G-VEVAL [32]. They usually support both reference-free and reference-based settings, where the former compares the candidate caption directly to the video without reference captions, while the latter uses both visual and textual references. We report results for both settings for completeness, but our primary focus is on the reference-free setting, which is more practical for real-world, in-the-wild videos. iii) Video-based metrics: These methods employ a video encoder to incorporate full video sequences as references. To the best of our knowledge, no existing metric in the literature falls into this category. Therefore, we adapt CLIPScore [9] to the recent advent of ViCLIP [35] as a stronger baseline, ViCLIPScore. Additionally, we compare VC-Inspector against the base model it builds upon, the vanilla Qwen2.5-VL model, to highlight the benefit of fine-tuning for evaluative reasoning.
Metrics.
In the caption evaluation literature, the effectiveness of evaluation metrics is typically assessed by measuring their correlation to human judgments. Following prior work [31, 32], we report both Kendall’s correlation ($\tau_{b}$) and Spearman’s rank correlation ($\rho$).
Implementation details.
VC-Inspector is developed for two model sizes, 3B and 7B, initialized from their corresponding Qwen2.5-VL pretrained weights. In all experiments, we train the model on 4 NVIDIA-A100 GPUs with a global batch size of 128 and a learning rate of 1e-4. We set both alpha and rank to 32 for the low-rank adaptation with a dropout rate of 0.05. During inference, we use a temperature of 0.0 for reproducibility.
5.2 Generalization of Quality Estimation
In this section, we evaluate the consistency of quality estimates using synthetic data, and validate the effectiveness of VC-Inspector by measuring its alignment with human judgments on the video caption evaluation dataset, VATEX-Eval. Given the limited availability of video caption datasets, we treat images as single-frame videos and extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF, to assess the generalization.
Evaluation with the synthetic data.
@@ 5.2 Generalization of Quality Estimation
EMScore [31] 28.94 40.77 20.21 29.24
CLIPScore [9] 28.10 39.65 18.00 26.14
Qwen2.5-VL-3B [2] 37.91 47.80 37.16 47.17
@@
Table 2: Correlation scores on the synthetic ActivityNet-FG-Eval [4] and YouCook2-FG-Eval [14] datasets. The best score is bolded.
Table 2 presents the correlation between metric scores and ground truth annotations in the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, both of which contain pseudo captions with varying degrees of factual inaccuracies. VC-Inspector, specifically finetuned for factual grounding, consistently outperforms baselines in differentiating incorrect captions. Notably, although it is only trained on ActivityNet-FG-It, its quality estimation generalizes effectively across visual domains. This suggests that the proposed data generation pipeline produces stable and reliable quality scores, rather than noisy outputs that fail to converge.
Evaluation on VATEX-Eval.
Table 5.2 presents our primary results: human correlation scores on the VATEX-EVAL [31] dataset. Baselines are grouped by reference type as outlined in section 5.1, and the three column-sections correspond to the No Reference, 1-Reference, and 9-Reference settings, respectively. Language-based metrics, which rely heavily on textual references, are not applicable to our target No Reference setting. Therefore, we focus our comparisons on multimodal methods that incorporate visual inputs. VC-Inspector consistently outperforms all evaluated metrics in the reference-free setting, particularly those based on CLIP embeddings, including CLIPScore, EMScore, FactVC, and PAC-S. When adapting CLIPScore to the recently introduced ViCLIP [35] (which adopts a video encoder), we observe a considerable gain over those image-CLIP-based approaches. Despite ViCLIPScore being a stronger baseline, it still underperforms relatively to VC-Inspector and its underlying model, as the context length of the ViCLIP model (i.e., 32) is much shorter than the 32K context length of ours, which supports a more flexible and comprehensive caption evaluation. We further compare VC-Inspector with G-VEval [32], a recent LMM-based evaluation method. However, G-VEval is based on GPT-4o, a proprietary, paid model with an estimated 200B parameters. In contrast, VC-Inspector is open-source, lightweight, and reproducible, with configurations available at 3B and 7B parameters depending on system requirements. While having a compact size, our 7B model achieves the highest correlation to human evaluations and could potentially enable on-the-fly quality estimation during training, making it a viable reward model in Reinforcement Learning (RL) applications. Although our primary focus is to compare with the reference-free metrics, it is noteworthy that VC-Inspector even surpasses the performance of most reference-based metrics.
Evaluation on Flickr8K.
@@
[…]
BLEU_1 [26] - - 12.65 16.52 28.70 36.88
BLEU_4 [26] - - 12.44 36.28 22.76 29.56
ROUGE-L [19] - - 12.94 16.89 23.94 31.06
METEOR [3] - - 16.68 21.80 27.64 35.76
CIDEr [33] - - 17.62 23.02 27.92 36.18
BERTScore [39] - - 15.24 19.82 25.05 32.37
SPICE [1] - - 14.80 18.78 27.41 35.40
SPICE-factual [18] - - 13.59 17.04 26.05 35.58
Soft-SPICE [18] - - 21.25 27.61 36.31 46.41
CLAIR [6] - - 36.00 - 34.80 -
[…]
CLIPScore [9] 22.33 29.09 27.39 35.49 35.21 45.28
EMScore [31] 22.88 29.79 28.63 37.05 36.66 47.00
FactVC [21] 22.79 29.69 28.78 37.22 36.18 46.33
PAC-S [30] 25.10 - 32.60 - 31.40 -
G-VEval [32] 39.40 - 44.90 - 48.10 -
[…]
ViCLIPScore [9, 35] 30.92 39.86 - - - -
Qwen2.5-VL-3B [2] 31.29 36.43 - - - -
Qwen2.5-VL-7B [2] 34.70 39.40 - - - -
@@
Table 3: Human correlation scores on the VATEX-EVAL [31] dataset. indicates results reported from [32] due to GPT-4 licensing issue. The best score for each column section is bolded. Please note that this work focuses on the No Reference setting. Our best model has outperformed all other models in this setting, while remaining competitive with metrics that rely on references.
@@
[…]
BLEU_1 [26] 32.20 17.90
BLEU_4 [26] 30.60 16.90
ROUGE [19] 32.10 19.90
METEOR [3] 41.50 22.20
CIDEr [33] 43.60 24.60
SPICE [1] 51.70 24.40
BERTScore [39] - 22.80
CLIPScore [9] 52.60 36.40
PAC-S [30] 55.50 37.60
[…]
CLIPScore [9] 51.10 34.40
PAC-S [30] 53.90 36.00
@@
Table 4: Correlation score ($\tau_{b}$) with human judgments on Flickr8k-Expert and Flickr8k-CF [10] dataset. The overall best scores are bolded and best of each section is underlined. Our model has outperformed even the reference-based methods.
To expand our evaluation beyond the limited availability of video caption datasets, we treat images as an extreme case of short videos with a single frame. Table 5.2 reports the human correlation score on two widely adopted benchmarks, Flickr8K-Expert and Flickr8K-CF. The former requires a more fine-grained differentiation among the captions, with ratings distributed across 4 levels, whereas the latter employs binary judgments. VC-Inspector is instruction-tuned with captions that exhibit varying degrees of factual inaccuracies, enabling nuanced evaluation of caption quality. This allows the model to perform effectively across both benchmarks, despite their differing rating scheme. In these evaluations, VC-Inspector remains the best-performing method under the reference-free setting, substantially narrowing the gap to inter-human correlation ($\tau_{b}\approx 73$[1], and even outperforming several reference-based metrics. These results demonstrate the strong generalization capability of VC-Inspector across visual domains and video lengths.
5.3 Ablation and Analysis
We conduct ablation studies on VC-Inspector components on the VATEX-Eval benchmark:
Data synthesis strategies.
Since both objects and actions are informative factual elements for a video, grounding the evaluator in factual understanding requires instructing the model to identify errors in these components. To this end, during the data generation process (outlined in section 4.1), we systematically altered both elements in ground truth captions to create pseudo captions for training. In this section, we ablate the impact of modifying these elements in the ActivityNet training set by evaluating three variants: i) Changing objects only, ii) Changing actions only, and iii) Changing both. As shown in Table 5, all variants achieve strong alignment with human ratings compared to EMScore. However, the variant that alters both objects and actions yields the best performance. These results highlight the importance of both factual elements, object and action, in capturing the context of video content for caption evaluation. They also demonstrate the robustness and generalization of the proposed paradigm across different factual errors.
Role of explanations.
Figure 4: Visual examples from ActivityNet-FG-Eval (top) and VATEX-Eval (others). VC-Inspector produces quality assessments highly consistent with ground truth scores, along with explanatory insights into factual errors (highlighted in red).
@@ 5.3 Ablation and Analysis
Table 5: Ablation study on synthetic data generation strategies. Human correlation scores ($\tau_{b}$, $\rho$) are reported on the VATEX-Eval.
During our data generation process, we systematically synthesize pseudo captions along with their corresponding quality estimates. This process yields an informative “side product": explanations that indicate where factual errors are located within the candidate captions. We leverage these explanations as an auxiliary supervision signal during training, enabling the model to implicitly learn factual grounding, which has been shown to be effective compared to the variant without explanations in Table 6. Not only are explanations useful in training, but they also enhance the interpretability of the model-based metric, providing transparency of why the scores are assigned to given captions. Further evaluation of the explanation quality is provided in the supplementary.
Metric stability.
To evaluate the stability of our metric, following prior work [36], we computed the Pearson correlation between scores obtained across two runs. We find a perfect correlation of 1.0 (as expected with temperature 0.0), confirming that VC-Inspector produces stable quality estimation across multiple runs.
@@
Computational efficiency.
We assess the computational efficiency by comparing the average runtime per video clip. EMScore [31], ViCLIPScore [35], and VC-Inspector require 0.42, 0.34, and 0.30 seconds per clip, respectively, on a single A100 GPU. VC-Inspector demonstrates superior efficiency compared to existing methods.
5.4 Qualitative Results
Figure 4 presents visual examples of VC-Inspector outputs on the ActivityNet-FG-Eval (top) and VATEX-Eval (middle, bottom) datasets. Without relying on reference captions, the model evaluates candidate captions based on the associated video content and produces quality scores that closely mirror human judgments. The explanation pinpoints the factual inaccuracies, such as incorrect objects and/or actions in the candidate captions, and assigns factually grounded scores accordingly. Additional qualitative results are provided in the supplementary.
6 Conclusion and Discussion

### v2

@@ 3.3 Training of VC-Inspector
The explanation $E$ is formatted in natural language using information collected during data generation (i.e., the list of changed objects and actions), as detailed in App.˜D. This can serve as extra supervision for the model to learn factual grounding and provide interpretable reasoning for this model-based evaluator at test time. Fig.˜4 presents an example of the VC-Inspector output. During training, the model is optimized with the standard language modeling loss Liu et al. (2023a).
Figure 4: Visual example on VATEX-Eval. VC-Inspector produces quality assessments consistent with ground truth scores, and factual error explanations (highlighted in red). More examples are in App.˜A.
@@ 4.1 Experimental Settings
We train VC-Inspector with the proposed instruction tuning dataset, ActivityNet-FG-It, which comprises 44K video-caption pairs, along with their quality scores and explanations. To evaluate whether VC-Inspector can generalize to other videos and visual domains, we further follow the same data generation pipeline in section 3.2 to create two evaluation datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, according to the ActivityNet Krishna et al. (2017) test set and YouCook2 Zhou et al. (2018) validation set. However, these are not truly labeled by humans. To evaluate whether the quality scores given by VC-Inspector are aligned with human evaluators, we utilized the widely adopted VATEX-EVAL dataset, specifically designed for evaluating video caption quality. It contains 6 captions with varying levels of quality per video, each rated by three human evaluators on a scale of 1 to 5. Since some videos become unavailable, we collect the remaining subset of 2,590 videos and the corresponding 15,540 candidate captions. Unless otherwise specified, all experiments were evaluated on the same dataset to ensure a fair comparison. We further extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF Hodosh et al. (2013), by viewing images as single-frame videos to test the generalization of the proposed metric. The former contains 17K image-caption pairs rated by three human experts on a scale of 1 to 4, while the latter collects binary quality assessments for 48K image-caption pairs from crowd sources.
Video preprocessing.
Each video is uniformly sampled into 32 frames and resized to a resolution of $224\times 224$ for both training and testing. For any image dataset, we treat them as one-frame videos.
@@
Baselines are organized into three categories: i) Language-based metrics solely rely on text reference without considering the visual input. Representative metrics are BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), CIDEr Vedantam et al. (2015), BERTScore Zhang et al. (2020), SPICE Anderson et al. (2016), SPICE-Factual Li et al. (2023), Soft-SPICE  Li et al. (2023) and CLAIR Chan et al. (2023b). ii) Image-augmented metrics incorporate images as references (alongside reference captions), for better capturing the semantic alignment between the visual input and the candidate caption, e.g., CLIPScore Hessel et al. (2022), EMScore Shi et al. (2022), PAC-S Sarto et al. (2023), FactVC Liu and Wan (2023), and G-VEVAL Tong et al. (2025). They usually support both reference-free and reference-based settings. We report results for both settings for completeness, but our primary focus is on the reference-free setting, which is more practical for real-world, in-the-wild videos. iii) Video-based metrics employ a video encoder to incorporate full video sequences. To the best of our knowledge, no existing metric in the literature falls into this category. Therefore, we adapt CLIPScore Hessel et al. (2022) to the recent advent of ViCLIP Wang et al. (2024) as a stronger baseline, ViCLIPScore. Additionally, we compare VC-Inspector against the base model it builds upon, the vanilla Qwen2.5-VL model, to highlight the benefit of fine-tuning for evaluative reasoning.
@@
EMScore Shi et al. (2022) 28.94 40.77 20.21 29.24
CLIPScore Hessel et al. (2022) 28.10 39.65 18.00 26.14
Qwen2.5-VL-3B Bai et al. (2025) 37.91 47.80 37.16 47.17
@@
Table 1: Correlation scores on the synthetic ActivityNet-FG-Eval Caba Heilbron et al. (2015) and YouCook2-FG-Eval Krishna et al. (2017) datasets. The best score is bolded.
Metrics.
In the caption evaluation literature, the effectiveness of evaluation metrics is typically assessed by measuring their correlation to ground truth. Following prior work Shi et al. (2022); Tong et al. (2025), we report both Kendall’s correlation ($\tau_{b}$) and Spearman’s rank correlation ($\rho$).
Implementation details.
VC-Inspector is developed for two model sizes, 3B and 7B, initialized from their corresponding Qwen2.5-VL pretrained weights. In all experiments, we train the model on 4 NVIDIA-A100 GPUs with a global batch size of 128 and a learning rate of 1e-4. We set both alpha and rank to 32 for the low-rank adaptation with a dropout rate of 0.05. During inference, we use a temperature of 0.0 for reproducibility. The computational analysis and the training hyperparameters are provided in App.˜B and App.˜C, respectively.
4.2 Generalization of Quality Estimation
Evaluation with synthetic data.
@@ 4.2 Generalization of Quality Estimation
Language-based
BLEU_1 Papineni et al. (2002) - - 12.65 16.52 28.70 36.88
BLEU_4 Papineni et al. (2002) - - 12.44 36.28 22.76 29.56
ROUGE-L Lin (2004) - - 12.94 16.89 23.94 31.06
METEOR Banerjee and Lavie (2005) - - 16.68 21.80 27.64 35.76
CIDEr Vedantam et al. (2015) - - 17.62 23.02 27.92 36.18
BERTScore Zhang et al. (2020) - - 15.24 19.82 25.05 32.37
SPICE Anderson et al. (2016) - - 14.80 18.78 27.41 35.40
SPICE-factual Li et al. (2023) - - 13.59 17.04 26.05 35.58
Soft-SPICE Li et al. (2023) - - 21.25 27.61 36.31 46.41
CLAIR Chan et al. (2023b) - - 36.00 - 34.80 -
Multimodal - image-based
CLIPScore Hessel et al. (2022) 22.33 29.09 27.39 35.49 35.21 45.28
EMScore Shi et al. (2022) 22.88 29.79 28.63 37.05 36.66 47.00
FactVC Liu and Wan (2023) 22.79 29.69 28.78 37.22 36.18 46.33
PAC-S Sarto et al. (2023) 25.10 - 32.60 - 31.40 -
G-VEval Tong et al. (2025) 39.40 - 44.90 - 48.10 -
Multimodal - video-based
ViCLIPScore Wang et al. (2024) 30.92 39.86 - - - -
Qwen2.5-VL-3B Bai et al. (2025) 31.29 36.43 - - - -
Qwen2.5-VL-7B Bai et al. (2025) 34.70 39.40 - - - -
@@
Table 2: Human correlation scores on the VATEX-EVAL Shi et al. (2022) dataset. indicates results reported from the original paper. The best of each column section is bolded. Please note that this work focuses on the No Reference setting. Our best model outperforms all other models in this setting, while remaining competitive with reference-based metrics.
Table˜1 presents the correlation between metric scores and ground truth annotations in the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, both of which contain pseudo captions with varying degrees of factual inaccuracies. VC-Inspector, specifically finetuned for factual grounding, consistently outperforms baselines in differentiating incorrect captions. Notably, although it is only trained on ActivityNet-FG-It, its quality estimation generalizes effectively across visual domains. This suggests that the proposed data generation pipeline produces stable and reliable quality scores, rather than noisy outputs.
@@
Reference-based
BLEU_1 Papineni et al. (2002) 32.20 17.90
BLEU_4 Papineni et al. (2002) 30.60 16.90
ROUGE Lin (2004) 32.10 19.90
METEOR Banerjee and Lavie (2005) 41.50 22.20
CIDEr Vedantam et al. (2015) 43.60 24.60
SPICE Anderson et al. (2016) 51.70 24.40
BERTScore Zhang et al. (2020) - 22.80
CLIPScore Hessel et al. (2022) 52.60 36.40
PAC-S Sarto et al. (2023) 55.50 37.60
Reference-free
CLIPScore Hessel et al. (2022) 51.10 34.40
PAC-S Sarto et al. (2023) 53.90 36.00
@@
Table 3: Correlation score ($\tau_{b}$) with human judgments on Flickr8k-Expert and Flickr8k-CF Hodosh et al. (2013) dataset. The overall best scores are bolded and best of each section is underlined. Our model outperforms even the reference-based methods.
Evaluation on VATEX-Eval.
Table 2 presents our primary results: human correlation scores on the VATEX-EVAL Shi et al. (2022) dataset. Baselines are grouped by reference type as outlined in section 4.1, and the three column-sections correspond to the No Reference, 1-Reference, and 9-Reference settings, respectively. Language-based metrics, which rely heavily on textual references, are not applicable to our target, No Reference setting. Therefore, we focus our comparisons on multimodal methods that incorporate visual inputs. VC-Inspector consistently outperforms all evaluated metrics in the reference-free setting, particularly those based on (image-based) CLIP embeddings. When adapting CLIPScore to the recently introduced ViCLIP Wang et al. (2024) (which adopts a video encoder), we observe a considerable gain over those image-CLIP-based approaches. Despite ViCLIPScore being a stronger baseline, it still underperforms relative to VC-Inspector and its underlying model, as the context length of the ViCLIP model (i.e., 32) is much shorter than the 32K context length of ours, which supports a more flexible and comprehensive caption evaluation. Furthermore, VC-Inspector outperforms G-VEval (Tong et al., 2025) despite being based on GPT-4o, a substantially larger proprietary model. This demonstrates that explicit factual grounding and explanation supervision are more critical than model scale alone. In contrast, VC-Inspector is open-source, lightweight, and reproducible, with configurations available at 3B and 7B parameters depending on system requirements. While having a compact size, our 7B model achieves the highest correlation to human evaluations and could potentially enable on-the-fly quality estimation during training, making it a viable reward model in Reinforcement Learning (RL) applications. Although we focus on the reference-free setting, it is noteworthy that VC-Inspector even surpasses the performance of most reference-based metrics.
Evaluation on Flickr8K.
To expand our evaluation beyond the limited availability of video caption datasets, we treat images as single-frame videos. Table 3 reports the human correlation score on two widely adopted benchmarks, Flickr8K-Expert and Flickr8K-CF. The former requires a more fine-grained differentiation among the captions, with ratings distributed across 4 levels, whereas the latter employs binary judgments. VC-Inspector is instruction-tuned with captions that exhibit varying degrees of factual inaccuracies, enabling nuanced evaluation of caption quality. This allows the model to perform effectively across both benchmarks, despite their differing rating scheme. In these evaluations, VC-Inspector remains the best-performing method under the reference-free setting, and even outperforming several reference-based metrics. These results demonstrate the strong generalization capability of VC-Inspector across visual domains and video lengths.
4.3 Ablation on Data Synthesis Strategies
@@
Table 4: Ablation study on synthetic data generation strategies. Human correlation scores ($\tau_{b}$, $\rho$) are reported on the VATEX-Eval benchmark.
Since both objects and actions are key factual elements in video captions, grounding the evaluator in factual understanding requires instructing the model to identify errors in these components. To this end, we systematically altered both elements in ground truth captions during data generation to create pseudo captions for training. In this section, we ablate the impact of modifying these elements in the ActivityNet training set by evaluating three variants: i) Changing objects only, ii) Changing actions only, and iii) Changing both. As shown in Table 4, all variants exhibit strong alignment with human ratings compared to EMScore Shi et al. (2022). However, the variant that alters both objects and actions yields the best performance. These results underscore the importance of both factual elements, objects and actions, for capturing video context in caption evaluation. They also demonstrate the robustness and generalization of the proposed paradigm across different factual errors.
4.4 Analysis of Explanations
Impact of explanation on performance.
Our data generation process produces not only pseudo captions and quality scores but also an informative “side product": explanations identifying factual errors in candidate captions. We leverage these explanations as auxiliary supervision during training, enabling the model to learn better factual grounding, which has been shown to be effective compared to the variant without explanations in Table 5. Beyond training benefits, explanations enhance the interpretability of the model-based metric, providing transparency of why specific scores are assigned.
Using explanations for caption refinement.
@@ 4.4 Analysis of Explanations
Having shown that explanations aids in evaluating caption quality, we now ask: can they also help improve the captions themselves? To assess this, we adopt Qwen2.5-VL-7B as the captioning model and use VC-Inspector-7B as the critique model that provides feedback during an iterative refinement process, repeated for up to 10 iterations. The prompt used for refinement is provided in App. D. We conduct this experiment with CAPability (Liu et al., 2025), a recent benchmark that evaluates caption quality across 12 distinct dimensions and two modalities (image and video), with human-annotated ground truth. Following its protocol, final captions are scored by GPT-4.1. Note that we report dimensions related to object semantics and actions in Table˜6, and provide the full results in Table˜8. We implement baselines using the same captioner but with different feedback strategies: (a) no feedback, and (b) feedback from the off-the-shelf Qwen2.5-VL-7B. Results show that the captioner guided by VC-Inspector consistently improves over both baselines on these dimensions, demonstrating the effectiveness of explanations for caption refinement. Notably, comparing the performance of (b) to (a), we find that undesired feedback to the captioner can even hurt the performance, significantly. This further underscores the importance of explicit factual grounding in VC-Inspector.
Evaluation of explanation quality.
For a more direct assessment of explanation quality, we randomly sample 100 captions from the VATEX-Eval dataset, and evaluate the explanations generated by VC-Inspector-7B, using 3 of the authors as annotators. For each explanation, annotators are provided with the corresponding video and candidate caption, and asked to rate the explanation on a Likert scale rather than binary judgments (e.g., good vs. bad), as explanations may be correct in some aspects while missing others. The explanations achieve a mean score of 4.33 with a standard deviation of 0.8. Inter-annotator agreement, measured by Krippendorff’s Alpha (ordinal), is 0.46, indicating moderate agreement. Notably, agreement within a margin of one point reaches 85.4%, suggesting strong consistency among annotators. In addition to human evaluation, we present language model based evaluation in App.˜B.
Feedback Provider
Task None Qwen2.5-VL-7B VC-Inspector-7B
Dynamic Object Num. 17.4 17.3 20.7
Action 58.2 18.8 58.6
Object Category 67.8 58.9 68.0
Object Number 28.8 20.4 29.3
Spatial Relation 56.6 45.7 57.1
Table 6: F1 scores across different tasks. Video-based tasks are highlighted in blue, while image-based tasks are highlighted in beige. Best results are shown in bold.
5 Conclusion